OpenStudy (anonymous):

The face of a tent is shaped like a triangle with an area of Fifteen feet squared. The area of the triangle is represented by the formula A equals one half b times h. The height of the tent is 4 feet. How wide is the base of the tent? 7.0 feet 7.5 feet 8.0 feet 8.5 feet

OpenStudy (anonymous):

7.5 ^.^

OpenStudy (anonymous):

Why is it 7.5?

OpenStudy (anonymous):

The formula is A = 1/2 * b * h Area= 15 height = 4 15 = 1/2 * b * 4 solve for b b = 15/2 = 7.5